OpenStudy (anonymous):

Real number situations... Sometimes,always or,never..... The sum of 2 rational numbers is an integer ???

OpenStudy (anonymous):

plllzz help

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110604132213AAUTCQN This link is from yahoo... Hope it helps.

OpenStudy (anonymous):

Therefore, I think I can conclude that this answer is always.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

No problem ;)

OpenStudy (anonymous):

oohh what would be a true problem i could use ? like a number problem adding two rational numbers to get a integer ??

OpenStudy (anonymous):

??????????????

OpenStudy (anonymous):

a/b + c/d = a*d/b*d + c*b/d*b = (ad + bc)/(bd) 4/1 + 6/2 = 4*2/1*2 + 6*1/2*1 = (8+6) / 2 = 14/2 = 7. Strike the 'Always'. The answer logically would be Sometimes. (because if you substituted the current '6' with '5', you get a decimal and decimals are not integers.)

OpenStudy (anonymous):

thanks again lol

OpenStudy (anonymous):

lol No problem. Sorry about the late reply lmao

OpenStudy (anonymous):

thats ok lol

OpenStudy (anonymous):

lol okay :D

OpenStudy (anonymous):

:) lmao

OpenStudy (anonymous):

;) haha

OpenStudy (anonymous):

:p lol jk

OpenStudy (anonymous):

haha. x] -.- Today's my first day of school and I already have homework, lmao.

OpenStudy (anonymous):

lol last tuesday was my first day and i had homework since then lol

OpenStudy (anonymous):

lol. You go to a private school? o.o

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

hmmm.... lol so you end school early?

OpenStudy (anonymous):

I guess idk lolz

OpenStudy (anonymous):

lol nice :D

OpenStudy (anonymous):

lol yepp wish it end very soon even though we just started..they need to quit hw uugg

OpenStudy (anonymous):

lol That's true. lmao. Hate homework.

OpenStudy (anonymous):

me to it sucks..stay school work.. i got stuff to do lol

OpenStudy (anonymous):

lol for sure haha

OpenStudy (anonymous):

lol yep

OpenStudy (anonymous):

;) lol

OpenStudy (anonymous):

:D lol